# The gauge symmetry - unbroken and broken - in the QCD ground state

### v2

@@ 7 The QCD full gluon propagator in the generalized gauge
The corresponding ST identity now becomes
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over(q^{2}+\lambda^{-1}\Delta^{2}_{t}(D))}.$ (7.6)
This is the generalized gauge since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. To our best knowledge the generalized gauge (7.6) has been derived for the first time in a such new manner. In the generalized gauge the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity. The functional dependence of the generalized gauge-fixing parameter $\xi$ (7.4) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.1), (7.5), and (7.6). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. 30 ; 31 ; 32 ; 33 . The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in 32 . This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (7.4) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (7.5) and the ST identity (7.6) for the regularized gluon fields.
The system of the regularized equations, consisting of the gluon SD eq. (7.1), the expression (7.5) and the ST identity (7.6), constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined. The tadpole term enters the full gluon self-energy linearly, see the expression (2.2). However, in the full gluon propagator (7.5) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (7.1). In the PT $q^{2}\rightarrow\infty$ limit, the term $(\Delta^{2}_{t}(D)/q^{2})$ in the expression (7.5) is to be suppressed, and it will be reduced to the PT QCD expression (5.10), reminding that $\xi_{0}=\lambda^{-1}$. The full gluon propagator will behave like the PT QCD gluon propagator at short distance ($q^{2}\rightarrow\infty$), namely $\sim 1/q^{2}$. This is the one of the necessary constraints that a theory is perturbative finite. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) 2 , evidently are beyond the scope of the present paper. Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry. Despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (7.5), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
8 Discussion
Let us now discuss some interesting features of the explicit presence of the tadpole term in the QCD ground state, i.e., at the fundamental quark-gluon level. First of all, let us underline once more that the tadpole term, existing in the QCD ground state, cannot be discarded in the QCD/YM theory by any means. This was exactly proven in section 6 and used in section 7. Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$. For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (7.1) up to the fist iteration term looks like
$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Bigl{[}\Pi(q^{2};D)+{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{]}-iL_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (8.1)
and it is worth reminding that the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms to one loop approximation have been omitted due to the relations (6.6)-(6.7). See appendix A as well, where this iteration is discussed in more detail. Let us also remind that the gluon eq. (7.1) is equivalent to the initial gluon SD eq. (2.1), shown in Fig. 1, and hence to eq. (2.5), only being re-written down in the form suitable for the NP renormalization program to be performed. The higher-order iteration terms will be much more singular with respect to the general ratio $(\Delta^{2}_{t}(D)/q^{2})$, since each iteration invokes additional powers of this ratio. Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms 34 , namely
$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},\tilde{\lambda}...)+O_{\mu\nu}(1/q^{2}).$ (8.2)
For simplicity, in this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. We have shown the singular part, associated with the transverse component of the full gluon propagator, since its longitudinal part is the regular function of the tadpole term, see expression (7.5). The dependence of the coefficients $\Phi_{k}(g^{2},\tilde{\lambda}...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above 23 . These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (8.2), which will remain massless, but depending on the tadpole term. Let us note that the same series can be obtained by expanding the initial NL transcendental relation (7.2) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
From the NL iteration expression (8.2) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 5. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (5.10) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluons will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions 26 , and thus is beyond the scope of the present work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
Eq. (7.5) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}_{g}$. The expression for the free massive vector particle propagator follows in a rather simple way, indeed. For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon pole mass $M^{2}_{g}$. In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[g_{\mu\nu}-(1-\lambda^{-1}){q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2}_{g})}\right],$ (8.3)
i.e., it will be expressed in the Stueckelberg gauge 4 . Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (7.6) by replacing there $q^{2}\rightarrow-q^{2}$. In other words, the Stueckelberg gauge (8.3) is a particular case of the generalized gauge (7.6). The PT QCD full gluon propagator (5.10) cannot generate the gluon pole mass at some finite point, since it remains always massless. The full gluon propagator (7.5) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields. This program will suppose to show that the full gluon propagator (7.5), and hence its free massive counterpart (8.3), can not be put on the mass-shell $q^{2}=M^{2}_{g}$. Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object. So that it will not appear in the physical spectrum (confinement of the massive gluon fields within our approach).
In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ 2 ; 35 ; 36 , which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT and full QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the asymptotic freedom (AF) regime 37 ; 38 .
Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime there is a scale breaking,
$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),\quad q^{2}\rightarrow\infty,$ (8.4)
which has been discovered many years ago 37 ; 38 . When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed. The expression (8.4) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$. $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD. However, in section 5, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT 2 ; 35 , namely
$M\sim\mu\exp(-1/b_{0}g^{2}),\quad\quad g^{2}\rightarrow 0,$ (8.5)
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discusses just above in this section. Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms).
9 Conclusions
Our formalism makes it possible to remove from the theory the QD constants of the gluon degrees of freedom in the same way as the QD constants of the quark degrees of freedom. This has lead to the discovery that in addition to the PT gluon field configurations in the QCD ground state (more precisely in its YM sector), preserving the exact gauge symmetry of its Lagrangian, there also exists such a type of the gluon field configurations which do not maintain this symmetry. But the PT renormalizable properties of the theory will not be affected. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq. (2.4) and Fig. 1. The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit 10 ; 23 ; 39 . By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand. It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig. 1), while all the other tadpole-like terms don’t. Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter 6 , indeed. Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way. In 10 the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one. However, from now on its existence cannot be undermined. Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum. Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
For the both types of the gluon field configurations have been derived the corresponding expressions for the full gluon propagators describing their propagation in the vacuum of QCD. For the second type the corresponding NP QCD full gluon propagator (7.5) is written down in the generalized gauge (7.6). From these expressions it is clearly seen that in the PT $q^{2}\rightarrow\infty$ limit eq. (7.5) coincides with eq. (5.10). This is an evidence that the PT renormalizable properties of the theory will not be affected by the new type of the gluon field configurations. In other words, the renormalizability of the theory will be maintained within our approach to the full gluon propagator. It is worth emphasizing once more that the massive and severely singular gluon field configurations can be described by eq. (7.5) only, while eq. (5.10) describes the propagation of the massless gluon modes with the PT IR singularities $\sim 1/q^{2}$ only. Our ground-breaking findings are as follows:
a. One of the transverse relations (3.5) is to be necessarily satisfied, in order to make the theory renormalizable in the PT sense.
b. The subtraction scheme has been formulated, which makes it possible to remove from theory all the quadratically UV divergent constants on the general basis.
c. The exact gauge symmetry of the YM Lagrangian is not a symmetry of its ground state. The dynamical source of this effect is the tadpole term. If it formally put zero then these gauge symmetries coincide.
d. The system of the relations (6.6)-(6.7) are exact mathematical results, i.e., they are not prescriptions.
e. The new NP approach to the YM theory and its ground state has been formulated, by keeping the corresponding ST identities always valid. It is based on the system of eqs.  (7.5)-(7.6), which is defined up to the corresponding invariant function with the known asymptotic properties. Eqs. (7.5)-(7.6) are suitable for the NP renormalization program to be performed.
f. The NL dependence $\xi$ on $\xi_{0}$ in the generalized gauge (7.4) in the PT $q^{2}\rightarrow\infty$ or, equivalently, formal $\Delta^{2}_{t}(D)=0$ limits becomes linear one $\xi=\xi_{0}$.
g. The DRM 6 ; 25 ; 26 ; 27 prescription to simply ignore all the QD constants (including the tadpole term) in the PT QCD has been put on a firm mathematical ground from now on, see eqs. (5.4)-(5.6).
In obtaining all these results no specific regularization schemes have been used. Neither special gauge choice nor any truncations, approximations and assumptions have been made either in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy. The exact analytic derivations have been done, based on a rigorous tensor algebra rules, such as the decomposition into the independent structures, the subtractions, etc.
10 Summary
It is shown in the most general and unique way that the gauge symmetry of the QCD Lagrangian is broken in its ground state, indeed. This has been achieved without solving the full gluon SD eq. (2.5), but recovering only its true dynamical and gauge structures in more detail in the framework of our formalism. So that, the significant peace of the information on these important issues was not missed within our general approach. Its breakthrough results can be summarized in terms of the following theorem:
Tadpole term existence and the true QCD ground state: Prove that in the QCD ground state exist such gluon field configurations which break/violet the gauge symmetry of its Lagrangian, but the PT renormalizability of the theory is maintained. Existence of this effect includes establishing the true role of the tadpole term in the QCD ground state.
Just this theorem has been proven above in this paper, which resolves the dilemma formulated at the end of section 2. In the proof of this theorem we have shown how to save the tadpole term in QCD and its ground state, without affecting the PT renormalizable properties of the theory, while removing all the other QD constants from it. The new derived solution with the gauge symmetry breaking will apparently satisfy the bounder conditions in QCD at the fundamental quark-gluon level: confinement at large distances ($q^{2}\rightarrow 0$) and AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 8. The solution with the gauge symmetry preservation can not satisfy them. All this underlines once more the importance of the theorem above.
The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the full gluon propagator (7.5), which is only regularized one at this stage. Unfortunately, this is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7. For this purpose it is necessary to take into account the transcendental and NL structure of eq. (7.2). They are also needed to be formulated in the most general and unique ways. To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well.
To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions. The confinement first and only after the PCAC phase transitions in QCD have to be understood 2 ; 23 . QCD is a self-consistent quantum gauge field theory. It does not need any extra degrees of freedom (such as Higgs fields 40 ) in order to generate a mass at the fundamental quark-gluon level.
Concluding, a few remarks are in order. The wide-spread opinion that the gauge symmetry cannot be broken/violated is misleading. The Elitzur’s theorem 41 has been proven in lattice field theory. It is supposed to be valid for the continuous theories as well, but if yes than only for the abelian systems, like QED. To the non-abelian systems, such as the YM fields, this theorem cannot be applied. In this paper we have just proven that the gauge symmetry of the YM theory is broken in its vacuum. The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there).
Acknowledgments
The authors are grateful to P. Levai, P. Forgács, J. Nyiri, T.S. Biró, Gy. Kluge, M. Vasúth for useful suggestions, remarks, discussions and help. The work was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) under the contract numbers OTKA K135515 and NKFIH 2019-2.1.11-TÉT-2019-00078, 2019-2.1.11-TÉT-2019-00050, the Wigner Scientific Computing Laboratory.
Appendix A The full gluon propagator up to one loop
Figure 2: The SD equation for the full gluon propagator up to one loop as present in 6 . It is useful to compare it with Fig. 1 and its description in the text after it.
In order to explicitly compare the tadpole term (2.4) with the other tadpole-type terms let us consider the iteration of the gluon SD eq. (2.1), shown in Fig. 1, up to one loop, shown in Fig. A.2. The last two terms in this figure appear because of the NL nature of the iteration series, when the next iterations contribute into the previous ones. That is the reason why such multiplied tadpole-type terms are not present in the skeleton loop decomposition for the full gluon propagator in Fig. 1, but have to be shown up in the iteration terms, seen in Fig. A.2. This iteration has been exactly calculated in 6 by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (5.5)-(5.6). It is worth emphasizing that in 6 this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$. It is well-known that in QCD the expansion in powers of the coupling constant does not make any sense because it is strong. The PT expansion in QCD makes sense only in the AF 37 ; 38 regime when it becomes weak. In order to perform analytical derivations below, we will take the interaction vertices as the point-like ones. The external gluon propagators will be taken as the free ones. All the skeleton loop propagators will remain as the full ones. Then the corresponding iteration will look like Fig. A.2, and hence can be called as quasi-one loop expansion, discussed in some details in section 8, when it was called as the cluster one. Note that below we do not show explicitly the dependence on the coupling constant, since our skeleton loop quark, gluon and ghost propagators are the full ones, so that the exact dependence on the coupling constant of the expressions below is not known. Also we will omit below all the colour group factors and other overall numerical numbers, since all of them are unimportant here.
However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig. A.2, for simplicity. From 6 by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq. (6.22) there, one finally obtains an analytical expression for this term as follows:
$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (A.1)
$N_{\rho\sigma}(l,q)=10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}+(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$ (A.2)
Putting in eqs. (A.1) and (A.2) $q=0$, one gets
$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (A.3)
Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtractions (4.1), as follows:
$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (A.4)

### v3

@@ Appendix A The full gluon propagator up to one loop
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (1.54)
see eq. (2.4) as well. Being the QD but already regularized constant, it is not connected to any external gluon momentum. In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure. Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state [6]. We have shown that in the PT QCD or,equivalently, particular QCD (described in section 6) it was a prescription to neglect it, but in the full QCD or, equivalently, NP QCD (investigated in sections 4 and 5) it should remain intact. At the time, all other tadpole-type terms have to be disregarded on the exact mathematical basis in both theories.
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (1.55)
where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only. Both gluon propagators in the t’ Hooft-Feynman gauge are $D^{0}_{\mu\mu^{\prime}}(p)=\delta_{\mu\mu^{\prime}}/p^{2}$ and $D_{\nu\zeta}(l)=\delta_{\nu\zeta}d(l^{2})/l^{2}$, where $d(l^{2})$ is the corresponding invariant function of the full gluon propagator. Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below. Using now the Euclidean space Feynman rules for the corresponding vertices present in [6], one arrives at
$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)=-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}}$ (1.56)
and
$T^{0}_{\nu\zeta\mu}(l,p)=2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (1.57)
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (1.58)
By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space [6] omitted) the first loop integral is zero, while the second one leads to
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (1.59)
where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared. It is one of the constants present in the relations (4.26). It should be discarded on the general mathematical basis within our approach, i.e., put zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit. Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq. (A.11) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.10), i.e., this pole is not a problem here. Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (1.60)
where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$. The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same. The altered rule [6] for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage. Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.
Appendix B The origin of the conventional QCD
In order to find the general solution to QCD, we present the full gluon self-energy (2.2) as a sum of the three independent terms, i.e., we distinguish between the gluon part and the tadpole term, though their sum forms the YM part of QCD, since the both are determined by the different self-interaction of multiple massless gluon modes. However, not doing so, one needs to present the full gluon self-energy in terms of the quark and YM compositions, i.e., one has to operate with
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D),$ (2.61)